40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1874     Entry time: Tue Aug 22 19:22:57 2017
Author: awade, Craig 
Type: DailyProgress 
Category: TempCtrl 
Subject: Fixing some of the lab temperature issues 

I removed some of the insulation around the north side of the tank to see beams going in when aligning the cavities.  When Craig and I replaced these, the stirring of air around the tank was enough to change the laser beat note by order 5 MHz over 15 -20 minutes.  

A rough calculation:

the absolute frequency of the laser is\nu = c/\lambda = 2.81 \times 10^{14}  Hz and expansion coefficient of fused silica is 5.5e-7 [1/K].  Frequency and length change are related by \Delta L /L = \Delta \nu/\nu.  So temperature is related to frequency shift by

\Delta T = \frac{\Delta \nu}{\alpha\nu} = \frac{\lambda}{c\alpha }\Delta \nu = 32 mK

This is the differential shift in cavity temperature over the 15-20 minute period.  At this rate of drift it is impossible to keep the PLL loop in range and means that our RefCavs are extremely sensitive to environmental drift

Some changes

- The intake of the hepa filters above the clean tent are directly below the output of the AC. There was some temporary aluminum foil put on the AC output vent to redirect air away from the top of the tent.  This was ripped and had holes about the edges so Craig is making a new one. We can make a more permanent thing out of Plexi glass later maybe. 

- We changed the room temperature from 72 F to 75 F to reduce the work the AC needs to do to maintain lab temperature (Michelson originally performed his experiments in a basement lab with no AC, they deliberately locked the door shut for >10 h stretches at time with no air circulation: they seem to have the right idea).  The real lab temperature measured with a temperature display was 19.9 C. We'll check again in the morning what this settles at. For now the BN has really changed a lot its moved from ~60 MHz down to 16 MHz.  This will take a while to settle.

 

ELOG V3.1.3-